OpenStudy (anonymous):

The speed at which a certain computer can access the Internet is 2 megabytes per second. How fast is this in megabytes per hour?

OpenStudy (anonymous):

60 seconds in a minute 2* 60 = 120 60 minutes in an hour 60 * 120 = 7200 in an hour

OpenStudy (aum):

1 hour = 60 minutes = 60 * 60 = 3600 seconds. Internet speed is 2 megabytes per second. That is, in one second 2 megabytes can be accessed. In 3600 seconds, 2 * 3600 = 7200 megabytes can be accessed. So the speed is 7200 megabytes per hour.

OpenStudy (anonymous):

Thank you all

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

To you that was probably really easy , but I just learned conversions today and got homework lol